The product of two numbers is 12. The sum of the larger number and twice the smaller number is 11. Find the two numbers.

Answer by jorel1380(3719)

Let m and n be the two numbers. Then:
2m+n=11
n=11-2m
m(11-2m)=12
11m-2mē=12
2mē-11m+12=0
(2m-3)(m-4)=0
m=3/2 or 4
Since 11-2(4)=3, then m must be 3/2, and n must be 8. ☺☺☺☺